OpenStudy (anonymous):

350 second song is divided into three parts; the beginning, middle, and the end. The beginning is one-sixth as long as the end. The middle part is one-half as long as the end. Find the length of each part.

OpenStudy (anonymous):

beginning=58.3 middle=175 end= 116.7

OpenStudy (anonymous):

You are amazing thank you!

OpenStudy (anonymous):

your welcome

OpenStudy (kropot72):

Let the length of the end = e seconds Length of begining = e/6 Length of middle = e/2 \[e+\frac{e}{6}+\frac{e}{2}=350\] \[\frac{6e+e+3e}{6}=350\] Now you can solve to find the length of the end (e) and calculate the begining and the middle.

OpenStudy (anonymous):

Thank you!

OpenStudy (kropot72):

You're welcome. Do you know how to solve for e ?

OpenStudy (anonymous):

Not exactly. Why don't you divide 350 by 6?

OpenStudy (kropot72):

You multiply both sides by 6 to get: \[\frac{10e}{6}\times 6=350\times 6\] \[10e=2100\]

OpenStudy (anonymous):

Ok I understand now thank you. I had the first part step correct but forgot to take the end product and divide it out with the original equation.

OpenStudy (kropot72):

Glad to know that you understand now :)